# Optimal Stopping with a Predicted Prior

### v1

@@ 7 Algorithm for MaxProb: Proof of Theorem 2
uniformly on compact subsets of $[0,1)$, which yields the above limit as $n\to\infty$. [^8] Rigorously, for any $\varepsilon>0$, uniform convergence over $0,1-\varepsilon]$ implies the convergence of $\int_{s}^{1-\varepsilon}\frac{\mathrm{e}^{-\frac{\gamma t}{1-s}}}{t}\,\mathrm{d}t$. The difference from the ideal term is bounded by $\int_{1-\varepsilon}^{1}\frac{1}{t}\,\mathrm{d}t=\mathcal{O}(\varepsilon)$, which vanishes as $\varepsilon\to 0$.
@@ 8.2 Representations of Algorithms
To derive the linear constraints in polytope $\mathcal{P}_{n,K,\tilde{F}}(\alpha,\beta)$, we consider two representations of online algorithms when the predicted prior is $\tilde{F}$ and the real prior is one of the conditional distributions in the above family $\mathcal{F}$. The first one relies on the conditional acceptance probabilities, which appeared in several previous works (e.g., buchbinderSecretaryProblemsLinear2014, correaSampledrivenOptimalStopping2021).
@@
Next, we show that we may without loss of generality focus on a class of minor-oblivious algorithms, whose decisions for whether accepting a best-so-far value $x_{t}$ at time $t$ is independent of smaller values $\bm{x}_{1:t-1}$ in the past.
@@
Recall that based on the realization of $\bm{x}_{1:t}$, the only information we may conclude about the prior $F_{k}$ is $k\geq\max\bm{x}_{1:t}$. Intuitively, there is no reason to make decisions based on the details of $\bm{x}_{1:t}$ other than a maximum $\max\bm{x}_{1:t}$. In this sense, minor-oblivious algorithms are without loss of generality. We formalize this characterization in the following proposition, which may be of independent interest. The proof is deferred to Section C.1.
@@
We artificially define $\textsc{Rej}_{0}(\ell)\coloneqq\mathds{1}{\scriptstyle{\{\ell=1\}}}$ to simplify some expressions in the rest of the section.
There is a one-to-one correspondence between the two representations (see Section C.2). We state below the direction to be used in the proof of Theorem 15.
@@ 8.3 Proof of Theorem 15
Given any minor-oblivious, $\alpha$-consistent, $\beta$-robust algorithm, we may assume without loss of generality that it rejects all values that are not best-so-far, since we consider MaxProb. We next verify that $y_{t,\ell}=\textsc{Rej}_{t}(\ell)$ is feasible w.r.t. polytope $\mathcal{P}_{n,K,\tilde{F}}(\alpha,\beta)$.
@@
$\textsc{Rej}_{0}(\ell)\,=\,1~,\qquad 0\,\leq\,\textsc{Rej}_{t}(\ell)\,\leq\,1~,\qquad\forall t\in[n],\;\forall\ell\in[K]~.$
@@
We did not explicitly write the optimality of $x_{t}$, i.e., $x_{t}=\max\bm{x}$, since $\textsc{Alg}=t$ implies $x_{t}=\max\bm{x}_{1:t}$ by our assumption without loss of generality that the algorithm rejects all values that are not best-so-far.
@@ 8.4 Numerical Results
$\lambda\cdot\alpha+(1-\lambda)\cdot\beta$
@@
We set $n=30$ and $K=1024$, and consider a predicted prior $\tilde{F}$ defined by $\tilde{f}(k)\propto\frac{1}{k}$ for $k\in[K]$. We solve the linear programs for $\lambda=\frac{i}{100}$ for $0\leq i\leq 100$ to get the curve shown in Figure 1(b). For details, see our code repository.
Appendix A Missing Proofs for MaxProb Consistency
A.1 Monotonicity of Consistency
@@
For any fixed $\lambda_{1},\lambda_{2}\in[0,1]$, the consistency $\Gamma_{n}\big(\theta^{*}_{n}\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}\big)$ is non-increasing in $n$.
@@
We show that for any $n$, it holds:
$\Gamma_{n+1}\big(\theta^{*}_{n+1}\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}\big)~\leq~\Gamma_{n}\big(\theta^{*}_{n}\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}\big)~.$
Consider a relaxed version of the problem with $n+1$ values, where the adversary tells the algorithm at the beginning the information about the smallest realized value, including its arrival time $\tau$ and its value $x_{\tau}$. We consider two policies.
In the first policy, we ignore the additional information and directly apply the algorithm parameterized by $\theta^{*}_{n+1}$. By definition, the consistency of this policy is $\Gamma_{n+1}\big(\theta^{*}_{n+1}\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}\big)$.
We now describe the second and optimal policy. Conditioned on $\tau$ and $x_{\tau}$, the remaining $n$ values are i.i.d. from $F$ restricted to $x_{\tau},+\infty)$, and their arrival times are i.i.d. uniform on $[0,1]$. Thus, by the optimality of $\theta^{*}_{n}$, any algorithm that rejects all values during $0,\lambda_{1}]$, accepts best-so-far values during $\lambda_{1},\lambda_{2}]$ based on a threshold function, and accepts best-so-far values during $\lambda_{2},1]$ has consistency at most $\Gamma_{n}\big(\theta^{*}_{n}\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}\big)$.
Since the second strategy is optimal, we get the stated inequality. ∎
A.2 Asymptotic Characterization of $\theta_{n}^{*}$
@@
For any fixed $s\in[0,1)$, we have
$\theta^{*}_{n}(s)~=~\frac{1}{1+\frac{\gamma}{(n-1)(1-s)}}+o\Big(\frac{1}{n}\Big)~,$
where $\gamma\approx 0.80435$ satisfies
$\sum_{k\geq 1}\frac{\gamma^{k}}{k!k}=1~.$

### v2

@@ 7.1 Asymptotic Performance Analysis
which holds uniformly on compact subsets of $[0,1)$. Substituting this limit into the integral yields the lemma. [^9] Rigorously, for any $\varepsilon>0$, uniform convergence over $0,1-\varepsilon]$ implies the convergence of $\int_{s}^{1-\varepsilon}\frac{\mathrm{e}^{-\frac{\gamma t}{1-s}}}{t}\,\mathrm{d}t$. The difference from the ideal term is bounded by $\int_{1-\varepsilon}^{1}\frac{1}{t}\,\mathrm{d}t=\mathcal{O}(\varepsilon)$, which vanishes as $\varepsilon\to 0$.
7.2 Tightness of the Trade-off
To complete the proof of Theorem 2, we show that no bi-criteria algorithm can achieve a better consistency-robustness trade-off.
First, we show that for any bi-criteria algorithm $\mathcal{A}(\tilde{F},\theta)$, the robustness guarantee given by Lemma 7 is tight.
Lemma 16.
For any $n\in\mathbb{N}$, any threshold function $\theta$, suppose $\lambda_{1}=\sup\{0\leq t\leq 1\colon\theta(t)=1\}$, $\lambda_{2}=\inf\{0\leq t\leq 1\colon\theta(t)=0\}$. Then, for any prior $F$ and any $\varepsilon>0$, there exists a predicted prior $\tilde{F}$ such that, when values are drawn i.i.d. from $F$, algorithm $\mathcal{A}(\tilde{F},\theta)$ accepts a maximum value with probability at most
$\min\{-\lambda_{1}\ln\lambda_{1},-\lambda_{2}\ln\lambda_{2}\}+\varepsilon~.$
Proof.
Fix any prior $F$ and $\varepsilon>0$. We prove two bounds, corresponding to “small” and “large” predicted priors.
Case 1: Small predicted prior.
Define $\tilde{F}_{\mathrm{small}}(x)=1$ for all $x\geq 0$ (equivalently, a point mass $1$ at 0). By monotonicity of $\theta$, for any $t>\lambda_{1}$ we have $\theta(t)<1$, so the second acceptance criterion $\tilde{F}_{\mathrm{small}}(x_{i})>\theta(t_{i})$ always holds. Hence, up to a measure-zero boundary at $t=\lambda_{1}$, the algorithm rejects all values before $\lambda_{1}$ and then accepts the first record, i.e., Dynkin’s continuous-time variant with transition point $\lambda_{1}$. By the standard analysis of that variant (e.g., Feldman et al. [27]), its MaxProb ratio equals $\lambda_{1}\ln\lambda_{1}$. Hence
$\operatorname{\mathbf{Pr}}\!\big[x_{\textsc{Alg}}(\tilde{F}_{\mathrm{small}},\theta)=x_{\textsc{Opt}}\big]\leq-\lambda_{1}\ln\lambda_{1}~.$
Case 2: Large predicted prior.
Since $F(\infty)=1$, there exists $m\geq 0$ such that
$F(m)^{n}\geq 1-\varepsilon~.$
Define
$\tilde{F}_{\mathrm{large}}(x)=\begin{cases}0,&x<m+1~;\\ 1,&x\geq m+1~.\end{cases}$
Let
$E\coloneqq\{X_{1}\leq m,X_{2}\leq m,\ldots,X_{n}\leq m\}~,$
so $\operatorname{\mathbf{Pr}}[E]\geq 1-\varepsilon$. On event $E$, all realized samples satisfy $\tilde{F}_{\mathrm{large}}(X_{i})=0$. Therefore, for any $t<\lambda_{2}$, since $\theta(t)>0$, the second acceptance criterion fails and the algorithm cannot accept before $\lambda_{2}$. Consequently, conditioned on $E$, this policy cannot outperform the policy that starts accepting records at time $\lambda_{2}$ (Dynkin variant), so
$\operatorname{\mathbf{Pr}}\!\big[x_{\textsc{Alg}}(\tilde{F}_{\mathrm{large}},\theta)=x_{\textsc{Opt}}\,\big|\,E\big]\leq-\lambda_{2}\ln\lambda_{2}~.$
Consequently,
$\operatorname{\mathbf{Pr}}\!\big[x_{\textsc{Alg}}(\tilde{F}_{\mathrm{large}},\theta)=x_{\textsc{Opt}}\big]\leq\operatorname{\mathbf{Pr}}[E]\cdot(-\lambda_{2}\ln\lambda_{2})+\operatorname{\mathbf{Pr}}[E^{c}]\leq-\lambda_{2}\ln\lambda_{2}+\varepsilon~.$
Finally, choose $\tilde{F}_{\mathrm{small}}$ if $\lambda_{1}\ln\lambda_{1}\leq-\lambda_{2}\ln\lambda_{2}$, and choose $\tilde{F}_{\mathrm{large}}$ otherwise. In both cases,
$\operatorname{\mathbf{Pr}}[x_{\textsc{Alg}}=x_{\textsc{Opt}}]\leq\min\{-\lambda_{1}\ln\lambda_{1},-\lambda_{2}\ln\lambda_{2}\}+\varepsilon~.$
If one insists on full-support continuous predicted priors, the same conclusion follows up to an arbitrarily small additive term by the smoothing argument in Section 2. ∎
We now complete the proof of Theorem 2. Fix any bi-criteria algorithm $\mathcal{A}(\tilde{F},\theta)$, and define
$\lambda_{1}^{\prime}=\sup\{0\leq t\leq 1\colon\theta(t)=1\}~,\qquad\lambda_{2}^{\prime}=\inf\{0\leq t\leq 1\colon\theta(t)=0\}~,$
$\beta\coloneqq\min\{-\lambda_{1}^{\prime}\ln\lambda_{1}^{\prime},-\lambda_{2}^{\prime}\ln\lambda_{2}^{\prime}\}~.$
By Lemma 16, for every $\varepsilon>0$ there exists a misspecification under which this algorithm succeeds with probability at most $\beta+\varepsilon$, hence its robustness is at most $\beta$.
Now let $\lambda_{1},\lambda_{2}$ be the two roots of $\lambda\ln\lambda=\beta$. By the concavity of $\lambda\ln\lambda$ over $[0,1]$, we have $\lambda_{1}^{\prime}\leq\lambda_{1}\leq\lambda_{2}\leq\lambda_{2}^{\prime}$. This means $\theta=\theta\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}$.
By Lemma 8, for these fixed $(\lambda_{1},\lambda_{2})$ and any $n$, the maximum consistency within the class $\mathcal{A}(\tilde{F},\theta\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}})$ is achieved by $\mathcal{A}(\tilde{F},\theta_{n}^{\scriptscriptstyle\mathrm{prob}}\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}})$. Therefore, no $\beta$-robust bi-criteria algorithm can have consistency larger than $\Gamma_{n}(\theta_{n}^{\scriptscriptstyle\mathrm{prob}}\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}})$. Taking $n\to\infty$ and applying Lemma 15, this upper bound converges to
$\alpha=\beta+\int_{\lambda_{1}}^{\lambda_{2}}\int_{s}^{1}\frac{\mathrm{e}^{-\frac{\gamma t}{1-s}}}{t}\,\mathrm{d}t\,\mathrm{d}s~.$
Hence, the trade-off in Theorem 2 is tight within the bi-criteria framework.
@@ 8.2 Representations of Algorithms
To derive the linear constraints in polytope $\mathcal{P}_{n,K,\tilde{F}}(\alpha,\beta)$, we consider two representations of online algorithms when the predicted prior is $\tilde{F}$ and the real prior is one of the conditional distributions in the above family $\mathcal{F}$. The first one relies on the conditional acceptance probabilities, which appeared in several previous works (e.g., Buchbinder et al. [8], Correa et al. [12]).
@@
Next, we show that we may without loss of generality focus on a class of minor-oblivious algorithms, whose decisions for whether accepting a record $x_{t}$ at time $t$ are independent of smaller values $\bm{x}_{1:t-1}$ in the past.
@@
Recall that based on the realization of $\bm{x}_{1:t}$, the only information we may conclude about the prior $F_{k}$ is $k\geq\max\bm{x}_{1:t}$. Intuitively, there is no reason to make decisions based on the details of $\bm{x}_{1:t}$ other than a maximum $\max\bm{x}_{1:t}$. In this sense, minor-oblivious algorithms are without loss of generality. We formalize this characterization in the following proposition, which may be of independent interest. The proof is deferred to Section D.1.
@@
We artificially define $\textsc{Rej}_{0}(\ell)\coloneqq 1$ to simplify some expressions in the rest of the section.
There is a one-to-one correspondence between the two representations (see Section D.2). We state below the direction to be used in the proof of Theorem 17.
@@ 8.3 Proof of Theorem 17
Given any minor-oblivious, $\alpha$-consistent, $\beta$-robust algorithm, we may assume without loss of generality that it rejects all values that are not a record, since we consider MaxProb. We next verify that $y_{t,\ell}=\textsc{Rej}_{t}(\ell)$ is feasible w.r.t. polytope $\mathcal{P}_{n,K,\tilde{F}}(\alpha,\beta)$.
@@
$\textsc{Rej}_{0}(\ell)\,=\,1~,\qquad 0\,\leq\,\textsc{Rej}_{t}(\ell)\,\leq\,1~,~~\forall t\in[n],\;\forall\ell\in[K]~.$
@@
We did not explicitly write the optimality of $x_{t}$, i.e., $x_{t}=\max\bm{x}$, since $\textsc{Alg}=t$ implies $x_{t}=\max\bm{x}_{1:t}$ by our assumption without loss of generality that the algorithm rejects all values that are not a record.
@@ 8.4 Numerical Results
$\lambda\cdot\alpha+(1-\lambda)\cdot\beta~,$
@@
We set $n=30$ and $K=1024$, and consider a predicted prior $\tilde{F}$ defined by $\tilde{f}(k)\propto\frac{1}{k}$ for $k\in[K]$. We solve the linear programs for $\lambda=\frac{i}{100}$ for $0\leq i\leq 100$ to get the curve shown in Figure 1(b). For details, see our code repository. [^10] https://github.com/billyldc/optimal-stopping-with-a-predicted-prior
Acknowledgement
We are grateful for the insightful discussions with Yiding Feng, Enze Sun, and Xiaowei Wu. We also wish to thank the participants of the Dagstuhl workshop 2025 “Online Algorithms Beyond Competitive Analysis” and the “Trends in Approximation and Online Algorithms (TAO)” workshop at FOCS 2025 for their valuable feedback.
Appendix A Derivation of Optimal Robust Threshold Functions
In this section, we prove Lemma 8. First, we derive the indifference (Bellman) equations for the bi-criteria algorithm under the MaxExp and MaxProb objectives. Then, we show that, for each objective, these equations uniquely determine the optimal threshold functions $\theta_{n}^{\scriptscriptstyle\mathrm{exp}}$ and $\theta_{n}^{\scriptscriptstyle\mathrm{prob}}$ when the predicted prior coincides with the true prior $F$. We first treat the MaxExp objective and then turn to MaxProb.
A.1 MaxExp: Optimal Threshold
Lemma 21 (Optimal threshold function for MaxExp).
For any $n\in\mathbb{N}$ and any prior $F$, the optimal threshold function $\theta_{n}^{\scriptscriptstyle\mathrm{exp}}$ that maximizes the consistency of $\mathcal{A}(F,\theta)$ among all $\theta$’s under MaxExp is uniquely determined at each time $t$ by $\theta_{n}^{\scriptscriptstyle\mathrm{exp}}(t)=F(x^{*}(t))$, where $x^{*}(t)$ is the unique solution of
$\phi(t,x)~=~0,$ (17)
with
$\phi(t,x)~\coloneqq~\operatorname{\mathbf{E}}_{y\sim F}[y\mid y\geq x]\left(1-\left(\frac{F(x)}{1-t+tF(x)}\right)^{n-1}\right)-x.$
Moreover, $\theta_{n}^{\scriptscriptstyle\mathrm{exp}}(t)$ is non-increasing in $t$ on $[0,1]$.
Proof.
Fix $t\in[0,1]$ and suppose the thresholds $\theta(s)$ have already been fixed for all $s\in(t,1]$. Let $x$ denote the value that arrives at time $t$, and assume $x$ is a record (else the algorithm cannot accept at $t$). By definition, the algorithm $\mathcal{A}(F,\theta_{n}^{\scriptscriptstyle\mathrm{exp}})$ accepts this value $x$ at time $t$ if and only if $F(x)>\theta_{n}^{\scriptscriptstyle\mathrm{exp}}(t)$. By optimality, the threshold chosen at time $t$ must make the expected rewards from accepting and rejecting this value $x$ equal.
Accepting. Accepting yields a reward $x$.
Rejecting. If the algorithm rejects $x$, then by the non-increasing threshold property, it must accept the first later value $y\geq x$. Thus, the continuation reward is
$\operatorname{\mathbf{E}}_{y\sim F}[y\mid y\geq x]\;\cdot\;\operatorname{\mathbf{Pr}}\!\big[\exists\,y\geq x\text{ after }t\mid x\text{ is a record at }t\big].$
To compute the existence probability, note that all other n1n-1 samples lie below $x$ with probability $F(x)^{n-1}$, and each sample is either below $x$ or arrives after $t$ with probability 1t+tF(x)1-t+tF(x). Conditioning on the latter event gives
$\operatorname{\mathbf{Pr}}\!\big[\exists\,y\geq x\text{ after }t\mid x\text{ is a record at }t\big]=1-\frac{F(x)^{n-1}}{(1-t+tF(x))^{\,n-1}}.$
This conditioning introduces the factor $(1-t+tF(x))^{n-1}$, because each of the other n1n-1 values must either arrive after time $t$ or fall below $x$. Equating the accept and reject rewards yields exactly the equation $\phi(t,x)=0$ in Equation 17.
By Lemma 22, for any continuous prior $F$ with full support on $0,\infty)$, any $n\in\mathbb{N}$, and any $t\in[0,1]$, the equation $\phi(t,x)=0$ has a unique solution $x^{*}(t)\in[0,\infty)$. Thus, defining $\theta_{n}^{\scriptscriptstyle\mathrm{exp}}(t)\coloneqq F(x^{*}(t))$ yields a unique optimal threshold at each time $t$.
Finally, by Lemma 23, $x^{*}(t)$ is non-increasing in $t$ on $[0,1]$. Since $F$ is non-decreasing, the same holds for $\theta_{n}^{\scriptscriptstyle\mathrm{exp}}(t)=F(x^{*}(t))$.
This completes the proof. ∎
We now establish existence, uniqueness, and monotonicity of the solution $x^{*}(t)$ to Equation 17. For the proofs it will be convenient to introduce some auxiliary notation. For $u\in[0,1)$ and $t\in[0,1)$, define
$g(u,t)\coloneqq\frac{1}{1-u}\left(1-\left(\frac{u}{1-t+tu}\right)^{n-1}\right),$
and for any $(t,x)$ with $t\in[0,1)$ and $x\geq 0$ define
$h(t,x)\coloneqq\frac{x}{g(F(x),t)}-\int_{x}^{\infty}y\,\,\mathrm{d}F(y).$
Using
$\operatorname{\mathbf{E}}_{y\sim F}[y\mid y\geq x]=\frac{\int_{x}^{\infty}y\,\,\mathrm{d}F(y)}{1-F(x)},$
one checks that for each fixed $t\in[0,1)$ and $x\geq 0$,
$\phi(t,x)=0\quad\Longleftrightarrow\quad h(t,x)=0.$ (18)
We first prove the existence and uniqueness of $x^{*}(t)$ for each $t$, and then show that $x^{*}(t)$ is non-increasing in $t$.
@@ A.1 MaxExp: Optimal Threshold
For any continuous prior $F$ with full support $0,\infty)$, any $n\in\mathbb{N}$, and any $t\in[0,1]$, Equation 17 has a unique solution $x^{*}(t)\in[0,\infty)$.
@@
When $t=1$, the factor
$1-\left(\frac{F(x)}{1-t+tF(x)}\right)^{n-1}$
equals 0 for all $x\geq 0$, so $\phi(1,x)=-x$ has a unique zero $x^{*}(1)=0$. We focus on the case $0\leq t<1$ below.
For fixed $t\in[0,1)$, the equivalence (18) implies that it suffices to show that $h(t,x)$ is strictly increasing in $x$ on $0,\infty)$ and crosses zero exactly once.
First, $g(F(0),t)=g(0,t)=1$, and by L’Hôpital’s rule
$\lim_{x\to\infty}g(F(x),t)=\lim_{u\to 1}g(u,t)=(n-1)(1-t).$
Hence
$h(t,0)$ $=-\int_{0}^{\infty}y\,\,\mathrm{d}F(y)\leq 0,$
$\lim_{x\to\infty}h(t,x)$ $=\lim_{x\to\infty}\frac{x}{(n-1)(1-t)}=\infty,$
so, by the Intermediate Value Theorem, $h(t,\cdot)$ has at least one root in $0,\infty)$.
We now show that $h(t,x)$ has at most one root in $0,\infty)$. It suffices to prove that $\partial_{x}h(t,x)>0$ for any $x\geq 0$. Differentiating in $x$, we obtain
$\partial_{x}h(t,x)=\frac{1}{g(F(x),t)}+\frac{x\,F^{\prime}(x)\big(g(F(x),t)^{2}-\partial_{u}g(F(x),t)\big)}{g(F(x),t)^{2}},$
where $F^{\prime}(x)$ denotes the density of $F$ (which exists almost everywhere). Since $g(u,t)>0$ for any $u\in[0,1)$, it remains to show that for any $u\in[0,1)$,
$g(u,t)^{2}-\partial_{u}g(u,t)\geq 0.$
Differentiating in $u$,
$g(u,t)^{2}-\partial_{u}g(u,t)=\frac{u^{n-2}\left[u^{n}-u(1-t+tu)^{n-1}+(n-1)(1-t)(1-u)(1-t+tu)^{n-2}\right]}{(1-u)^{2}(1-t+tu)^{2n-2}}.$
Let $z(u,t)$ denote the bracketed term in the numerator. Then $z(u,1)=0$ for any $u\in[0,1)$, and
$\partial_{t}z(u,t)=-(n-1)(1-u)^{2}(1-t+tu)^{n-3}\left[(1-t+tu)+(1-t)(n-2)\right]\leq 0.$
Thus $z(u,t)\geq z(u,1)=0$ for any $u,t\in[0,1)$, which implies $g(u,t)^{2}-\partial_{u}g(u,t)\geq 0$ and hence $\partial_{x}h(t,x)>0$ for all $x\geq 0$.
Therefore, for each fixed $t\in[0,1)$, the function $h(t,\cdot)$ is strictly increasing on $0,\infty)$ and crosses zero exactly once. By (18), the equation $h(t,x)=0$ has a unique solution $x^{*}(t)\in[0,\infty)$ for each fixed $t\in[0,1)$; and together with the case $t=1$, this proves the claim and the lemma. ∎
@@
For any continuous prior $F$ with full support $0,\infty)$ and any $n\in\mathbb{N}$, the unique solution $x^{*}(t)$ of Equation 17 is non-increasing in $t$ on $[0,1]$.